OpenStudy (anonymous):

Find the sum of the given arithmetic series. 26 + 19 + 12 + 5 + ... + (–37)

OpenStudy (anonymous):

since this is an arithmetic sequence and here first term is 26 and the common difference is -7 we need to know the number of terms n here

OpenStudy (anonymous):

so -37 = 26 +(n-1)*(-7) -63 =(n-1)*(-7) or n-1 =9 n=10

OpenStudy (anonymous):

sum of this AP can be found using sum = (no. of terms) * (first term +last term)/2